A rectangular field is 95 yards long and 65 yards wide.
Give the length and width of another rectangular field that has the same perimeter but a larger area.

Answer by rfer(16322)

95*65=6175 sq yds
P=320 yds
320/4=80
80^2=6400 sq yds